OpenStudy (anonymous):

derivative of arcsin(cos(2x))

OpenStudy (anonymous):

Let \[f(x)=\arcsin\left(\cos(2x)\right)\] So : \[f'(x)=-2\sin(2x)\times \frac{1}{\sqrt{1-\cos^2(2x)}}\]

OpenStudy (anonymous):

i would do something slightly different

OpenStudy (anonymous):

or just look more carefully at what you have in the answer above

OpenStudy (anonymous):

since \(1-\cos^2(2x)=\sin^2(2x)\) the answer given above is \[f'(x)=-2\sin(2x)\times \frac{1}{\sqrt{1-\cos^2(2x)}}\] \[-2\sin(2x)\times \frac{1}{\sqrt{\sin^2(2x)}}=-\frac{\sin(2x)}{|\sin(2x)}\]

OpenStudy (anonymous):

typo there, i forgot the 2 in the numerator the point is that the derivative is either 2 or -2 depending on the "sign" of \(\sin(2x)\)